a teacher purchased 69 tickets cost 9 $ for children 12 for adults she paid 678$ how many of each were sold
Answer by richwmiller(17219)

x+y=69,
9*x+12*y=678
x=69-y
9*(69-y)+12*y=678
621-9y+12*y=678
3*y=57
y=19
x=69-y
x=50 y=19
check
50+19=69
9*x+12*y=678
9*50+12*19=678
450+228=678
678=678
ok